OpenStudy (anonymous):

Also there is a math problem I need help on it is x squared times 9 cubed equals

OpenStudy (fibonaccichick666):

first, can you write it out in math terms?

OpenStudy (snowcrystal):

\[x^2 * 9^3\]

OpenStudy (anonymous):

[x ^{2} \times 9^{3} =?\]

OpenStudy (fibonaccichick666):

ok so, what does 9^3 =?

OpenStudy (anonymous):

9^3 = 9*9*9

OpenStudy (snowcrystal):

x^2 = x*x

OpenStudy (anonymous):

\[\Huge x ^{2} \times 9^{3} =?\]

OpenStudy (anonymous):

Forget it

OpenStudy (fibonaccichick666):

? why?

OpenStudy (anonymous):